OpenStudy (andrewthecookie):

What proportion should I make: 50% of 70?

OpenStudy (andrewthecookie):

@mathstudent55 @waterineyes @Zarkon @TheSmartOne

OpenStudy (andrewthecookie):

@SolomonZelman

OpenStudy (andrewthecookie):

@AriPotta

OpenStudy (godlovesme):

What's the total in percentage?

OpenStudy (aripotta):

50% of 70 is the same as 70/2!

OpenStudy (godlovesme):

50/100= x/70

OpenStudy (andrewthecookie):

It just says Which proportion can be used to solve the percent problem? What is 50% of 70

OpenStudy (anonymous):

50% of 70 means half of 70 divide 70 b y 2

OpenStudy (andrewthecookie):

ill send as a photo tell me if I'm correct

OpenStudy (aripotta):

oh i see what you mean. it should be \[\frac{ 50 }{ 100 }=\frac{ x }{ 70 }\]

OpenStudy (andrewthecookie):

oh... is there any correct answer there?

OpenStudy (andrewthecookie):

@Nnesha

OpenStudy (aripotta):

yes, the first choice

OpenStudy (andrewthecookie):

ok thx Ari!

OpenStudy (aripotta):

no problemo :)

OpenStudy (godlovesme):

Thanks for the medal @aripotta :)